OpenStudy (anonymous):

Solve the equation. Show your work. 8x – 12 = 4x + 24

OpenStudy (anonymous):

8x - 4x = 4x 24 + 12 = 36 36/4x = 9 x = 9

OpenStudy (anonymous):

To check: 8(9) = 72 72 - 12 = 60 4(9) = 36 36 + 24 = 60

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you have help me out so much i cant thank u enough

OpenStudy (anonymous):

Your welcome, let me know if you need any more help.

OpenStudy (anonymous):

ok thanks